OpenStudy (mendicant_bias):

Having trouble taking the laplace of t^2 times the unit step function, any help is appreciated.

OpenStudy (mendicant_bias):

I just don't know what steps I should take algebraically to get t^2 into the form of t-1; what i'm trying to do is take\[\mathcal{L}\left\{t^{2}\mathcal{U}(t-1)\right\}\]

OpenStudy (mendicant_bias):

Somehow, I need to change\[f(t)=t^2\]into the form\[f(t)=(t-1)^2\]

OpenStudy (anonymous):

ok

OpenStudy (mendicant_bias):

I'm thinking it might have something to do with\[(t-1)^{2}=t^2-2t+1\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

that is right

OpenStudy (anonymous):

ok.....45328

OpenStudy (mendicant_bias):

I'd really rather you not waste my time. @satellite73

OpenStudy (anonymous):

so that wod be 54 6578

OpenStudy (mendicant_bias):

Alright. Now that that's out of the way.

OpenStudy (mendicant_bias):

@Concentrationalizing , Any idea?

OpenStudy (mendicant_bias):

Wait a second, I think I got it.

OpenStudy (mendicant_bias):

\[(t-1)^2=t^2-2t+1; \ \ \ (t-1)^2+2t-1=t^2.\]Now, the other term involving t that has been generated needs to be dealt with(?)

OpenStudy (mendicant_bias):

Subtract an additional one from the laplace and take the laplace of it as well?

OpenStudy (anonymous):

yes

OpenStudy (mendicant_bias):

\[\mathcal{L}\left\{[(t-1)^2+2(t-1)]\mathcal{U}(t-1)\right\}+\mathcal{L}\left\{1\right\}\]@Concentrationalizing , does this look right?

OpenStudy (anonymous):

yes

OpenStudy (mendicant_bias):

@Hero (or any mod), take a look at this, at least so you have an idea what might be a problem in the future.

OpenStudy (anonymous):

do you what me to fan you and medal you?

OpenStudy (anonymous):